what is the slope of the line given the equation y+2=3(x-5)
Answer by stanbon(75887)

what is the slope of the line given the equation y+2=3(x-5)
----
Write the equation in slope-intercept form:
y+2 = 3x-15
y = 3x - 17
---
slope = 3
=================
Cheers,
Stan H.
=================